# JWST Discovery of a High-Redshift Tidal Disruption Event Candidate in COSMOS-Web

### v1

@@ Abstract
The rates and properties of tidal disruption events (TDEs) provide valuable insights into their host galaxy central stellar densities and the demographics of their central supermassive black holes (SMBHs). TDEs have been observed only at low redshifts ($z\lesssim 1$), due to the difficulty in conducting deep time-domain surveys. In this work, we present the discovery of a high-redshift TDE candidate, HZTDE-1, in the COSMOS-Web survey with JWST’s NIRCam, using a novel selection technique based on color and morphology. We first outline a methodology for identifying high-z TDEs in deep infrared imaging surveys, leveraging their unique spectral energy distributions (SEDs) and morphologies of these transients. We apply this technique to COSMOS-Web in filters F115W, F150W, F277W, and F [^4] W, and identify HZTDE-1, a transient point source relative to archival UltraVISTA infrared observations. If we assume it is a TDE, we estimate its photometric redshift to be $z=5.02^{+1.32}_{-1.11}$. HZTDE-1 cannot be explained by reasonable supernova or AGN models. However, we cannot rule out a superluminous supernova at $z\gtrsim 3$. If confirmed with follow-up observations, HZTDE-1 would represent the highest-redshift TDE discovery to date, and would suggest an enhancement of the TDE rate in the high-redshift universe. Our method, which can be applied to future deep surveys with JWST and Roman, offers a pathway to identify TDEs at $z>4$ and probe black hole demographics at early cosmic times.
@@ 1 Introduction
A tidal disruption event (TDE) occurs when a star passes close enough to a supermassive black hole (SMBH) that the tidal forces on the star overcome its self-gravity, and destroy the star (Hills, 1975; Rees, 1988; Evans & Kochanek, 1989; Ulmer, 1999). This process leads to a luminous flare ($\sim 10^{43}$erg/s), as approximately half of the mass of the star is accreted and the other half is ejected at high velocity (Strubbe & Quataert, 2009; Gezari et al., 2012). This highly transient accretion event contrasts with the more long term, sustained accretion that is associated with active galactic nuclei (AGN) (Begelman et al., 1984; Krolik, 1999; Event Horizon Telescope Collaboration et al., 2019). Despite decades of dedicated efforts, several outstanding questions surrounding the formation and evolution of SMBHs persist. In particular, the uncertain origin of the first SMBHs has outlined the need for robustly measuring the early Universe SMBH mass distribution (Volonteri et al., 2003; Begelman et al., 2006). TDEs are one of the only phenomena, alongside AGN, that present a luminous observable very close to a SMBH. Unlike AGN, which prefer very massive SMBH hosts (Kauffmann et al., 2003), TDEs of main-sequence stars by non-rotating (Schwarzschild) black holes can only occur for lower mass SMBHs because the disrupting black hole has an upper mass limit of $M_{BH}<10^{8}M_{\odot}$ (Hills, 1975). Beyond this mass, the tidal radius $R_{t}=R_{\star}(M_{BH}/M_{\star})$ is within the Schwarzschild radius $R_{s}=2GM/c^{2}$. The TDE rate’s sensitivity to the black hole mass make TDEs an excellent probe of the SMBH mass function (Stone & Metzger, 2016; Kochanek, 2016). The TDE rate is also affected by the stellar mass function, so a redshift-evolving TDE rate could also be sensitive to the initial mass function (IMF) (Magorrian & Tremaine, 1999). Additionally, stellar dynamics calculations imply that galaxies with lower-mass SMBHs have higher intrinsic TDE rates (Stone & Metzger, 2016). Therefore, TDEs are ideal probes of distant, lower-mass SMBHs and the stars which surround them.
@@
Typically TDEs are discovered through wide-field, time-domain optical surveys (e.g. the Zwicky Transient Facility (ZTF) Graham et al., 2019; Bellm et al., 2019). They are slow (lasting 100100100100200200200200 days), blue (peaking in the UV) transients that do not typically cool as they fade (Gezari, 2021). TDEs last on scales of months, typically fading from peak in $\sim 200$ days (van Velzen et al., 2021). The origin of the early-time UV/optical emission is debated, and still not fully understood (Ryu et al., 2023; Dai et al., 2018; Roth et al., 2020). In this work, we adopt a fully empirical, observation-driven model of TDE spectra and light curves.
@@
Previously, Kochanek (2016) has predicted that high-redshift TDEs would not be observable in ground-based wide-field time domain surveys. Here, we extend to high redshifts by leveraging deep space-based imaging without the use of time series. We assert in a companion work Karmen & et al. (in prep.) that, given the observed evolution of galaxy and SMBH properties with redshift, the TDE rate at high redshifts is likely greatly enhanced. As a result, in upcoming surveys it may be possible to study populations of TDEs at high redshift. Inayoshi et al. (2023) predicts the properties of high-redshift TDEs embedded within gas-rich AGN environments, in which the TDE occurrence rates are enhanced due to their compact and dense environments, compared to typical TDEs which occur in post-starburst galaxies. In this work, we complement this by a search exclusively for TDEs in galaxies not containing luminous AGN in order to utilize observed properties from low-redshift TDE searches, which exclude all AGN-hosting galaxies.
@@
In this work, we report the discovery of a $z=5.02^{+1.32}_{-1.11}$ TDE candidate in the COSMOS-Web survey (Casey et al., 2022). In Section 2, we present a new methodology for identifying TDE candidates in deep imaging using colors and morphology. We compare to potential contaminants, and pose methods of confirming candidate TDEs. In Section 3 we apply this methodology to the COSMOS-Web survey data, and use historic imaging of the COSMOS field to evaluate the potential variability of objects we select. We present the properties of HZTDE-1, a transient high-redshift TDE candidate in Section 4. We compare potential supernova models for this object to a TDE explanation. We discuss the need for spectroscopy of this object, and its implications if validated in Section 5. We summarize our work in Section 6.
@@
3
2 Identifying High-Redshift TDEs in Deep Surveys
In this section, we model the signatures of TDEs similar to the local population, and develop a method for identifying them in wide field infrared surveys using the colors and morphologies. To first order, the UV/optical SED of a TDE is well-approximated by a black body with a typical temperature of $\sim 10^{4.3\pm 0.15}~{}K$ (van Velzen et al., 2020, 2021; Hammerstein et al., 2023; Yao et al., 2023). We simulate TDEs with the the same temperature distribution as the latest sample of spectroscopically confirmed TDEs from ZTF (Yao et al., 2023), and simulate their volumetric density using the ZTF $g$-band luminosity function. TDEs typically do not cool as they fade, so we assume constant-temperature blackbodies for their entire evolution. Their light curves are observationally well-parameterized in van Velzen et al. (2021) as follows:
$L_{\nu}(t)$ $=L_{\nu_{0}\,\rm peak}~{}\frac{B_{\nu}(T_{0})}{B_{\nu_{0}}(T_{0})}$
×$\times\begin{cases}e^{-(t-t_{\rm peak})^{2}/2\sigma^{2}}&t\leq t_{\rm peak}\\ e^{-(t-t_{\rm peak})/\tau}&t>t_{\rm peak}\\ \end{cases}$ (1)
with the luminosity in the $g$-band in erg s-1 $44.6842.68<\log(L_{g})<$ 44.68, the Gaussian rise time $1.30.4<\log(\sigma)<$ 1.3 days and the exponential decay time $2.31.2<\log(\tau)<$ 2.3 days from ZTF. We allow TDEs to fade until they reach the magnitude limit of the filter/survey observing them, to account for the time evolution of the TDE.
@@ 2.1 The Host model
We base our predictions of properties of TDE host galaxies on local observations of TDE hosts. It has been observed that TDE hosts are disproportionately in the green valley with compact nuclei (Hammerstein et al., 2021), and may show a preference for post-starburst galaxies (French et al., 2016). Because of this, we base our host galaxy models on the total sample of 30 ZTF TDE hosts from (Hammerstein et al., 2021) to capture their observed properties. To extend the observed host spectra to their rest-frame UV, we use Prospector to model their spectra and perform synthetic photometry (Johnson et al., 2021). We use the parameters from the original fits of the ZTF TDE hosts in Hammerstein et al. (2021), which are estimated by maximum a posteriori fits of Flexible Stellar Population Synthesis models (Conroy et al., 2009) to broadband photometry from SDSS, Pan-STARRS, and GALEX. When calculating properties of TDE hosts, we simulate every known TDE host to account for their full diversity of properties. When simulating observed TDEs in a survey, we include the host spectrum by adding a mean template, derived from the Prospector models, to the TDE spectrum before performing synthetic photometry.
@@
We also use these ZTF TDE hosts as a basis for the physical sizes for typical TDE hosts. We query SDSS photometry for the TDE hosts in ZTF, and convert their angular sizes to physical sizes (Blanton et al., 2017; Hammerstein et al., 2023). We find that ZTF TDE host galaxies contain 90% of petrosian flux in a diameter of $8\leq D_{90}\leq 20$ kiloparsecs. This corresponds to range of angular diameters between $1.4$ and $4.7$ arcsec at $z=8$. This angular size will increase at even higher redshifts due to the decrease in angular diameter distance with increasing redshift beyond $z\sim 1.5$ (Hogg, 1999). As NIRCam has a resolution of 0.07 arcsec at 2 microns, all TDE host galaxies above the survey magnitude limit will be resolvable as an extended source in COSMOS-Web. At early epochs ($z\sim 7$), galaxies are expected to be intrinsically smaller and fainter (Grazian et al., 2012; Kawamata et al., 2018), such that even TDEs in unusual hosts will likely be point sources. Because of this, we predict that the majority of TDEs at high redshifts will outshine their hosts in the COSMOS-Web filters and be point sources.
@@
Because all of the host galaxies we observe are at $z<1$, the galaxies we use are of higher mass and more luminous than those which we expect at high redshifts. High redshift galaxies are expected to be, and now observed to be, primarily lower-mass, blue, star-forming galaxies. While there is a minority of these galaxies which are exceptionally luminous without SMBHs above the Hills mass due to extremely active star formation (e.g. Franco et al., 2024; Casey et al., 2024; Castellano et al., 2024), it would be extremely difficult to detect a TDE in one because its star formation is bright and blue, making the TDE emission a sub-dominant component of its UV SED. Therefore, we do not consider them in our analysis. This makes the local TDE hosts we observe effectively upper limits for higher-redshift host galaxy luminosity. In summary, TDEs trace lower-mass SMBHs which trace, in turn, lower mass galaxies. These are expected to be even less massive at high-redshifts, meaning we expect the majority of TDEs to be hostless.
@@ 2.3 High Redshift TDEs
We find that, given our simulated TDEs and assumptions about their typical host galaxies, most high-redshift ($z>4$) TDEs will be point sources with unique colors. At $z>4$, a typical TDE host galaxy drops in apparent magnitude below the detection limits of the COSMOS-Web survey, as seen in Figure 1. A TDE near peak brightness is visible to beyond $z\sim 10$. The SED of a high-redshift TDE is dominated by its blackbody emission, which is redshifted from the rest-frame UV to the observer-frame near-infrared NIR. Because a redshifted blackbody is another blackbody, there is a degeneracy in color-color space between temperature and redshift, meaning all idealized TDEs lie upon a nearly straight line in 2D color-color space (slightly curved due to relativistic Doppler shift and dust extinction). In our selection for high-redshift TDEs, we restrict to a region that follows the color-color relationship in any given color combination. This is shown as the three dark purple lines outlining the TDE region in Figure 3. We also fit bounds to each individual color, which, together with the parabola fits, are presented in Table 1. We fit a parabola to the color-color relationship of TDEs in each filter combination, parameterized as:
@@
where $C_{1}$ and $C_{2}$ are two colors. The values of coefficients $a_{1}$, $a_{2}$, and $a_{3}$ for each pair of colors are listed in Table 1, alongside the bounds on $C_{1}$ and $C_{2}$ form our simulations. Additionally, we fit the minimum apparent magnitude expected for a TDE at any redshift of a given color, shown in Figure 3, as
@@
To check our blackbody idealization, we take the following TDEs that have Hubble UV spectroscopy, and simulate them at high redshifts: iPTF16fnl, ASASSN-15oi, AT2022dsb, AT2018bsi, and ASASSN-14li (Charalampopoulos et al., 2022; Cenko et al., 2016; Blagorodnova et al., 2019; Brown et al., 2018; Hung et al., 2019; Hoogendam et al., 2024). We extend the UV spectrum into the optical ($>2600$) with a blackbody with the temperature fit by Mummery & van Velzen (2024), redshift the TDE spectra, and add in Lyman absorption shortward of rest-frame 1216Å (Inoue et al., 2014). We show these templates in Figure 2. These TDEs have colors centered on the relationship predicted, and scatter due to emission and absorption features, as well as continuum excess farther into the UV. This scatter, typically $\sim 0.2$ mag, defines the region near our idealized color-color relationships in which we expect actual TDEs to exist (the shaded region on the figure).
@@
Figure 2: Example templates used to test the TDE color-magnitude cuts, shifted to the rest frame. The red line is the HST spectrum, purple dotted is the blackbody used to extend the spectrum, and photometry is in blue. We switch from the HST spectrum to the blackbody at rest-frame wavelength $2600$, which is shown in the changing opacity of each line. We simulate this spectrum at higher redshifts to test our color cuts.

### v4

@@ Abstract
The rates and properties of tidal disruption events (TDEs) provide valuable insights into their host galaxy central stellar densities and the demographics of their central supermassive black holes (SMBHs). TDEs have been observed only at low redshifts ($z\lesssim 1$), due to the difficulty in conducting deep time-domain surveys. In this work, we present the discovery of a high-redshift TDE candidate, HZTDE-1, in the COSMOS-Web survey with JWST’s NIRCam, using a novel selection technique based on color and morphology. We outline a methodology for identifying high-z TDEs in deep infrared imaging surveys, leveraging their unique spectral energy distributions (SEDs) and morphologies of these transients. While focused on TDEs, this methodology could also be applied to find other UV-bright transients, such as superluminous supernovae (SLSNe). We apply this technique to COSMOS-Web in filters F115W, F150W, F277W, and F [^4] W, and identify HZTDE-1, a transient point source relative to archival UltraVISTA infrared observations. If we assume it is a TDE, we estimate a photometric redshift of $z=5.02^{+1.32}_{-1.11}$. HZTDE-1 cannot be explained by reasonable supernova or AGN models. However, a SLSN at $z\gtrsim 3$ can also plausibly explain this transient and would be the highest-redshift known SLSN. If confirmed with follow-up observations, HZTDE-1 would represent the highest-redshift TDE discovery to date, and suggest an enhancement of the TDE rate in the high-redshift universe. Our method, which can be applied to future deep surveys with JWST and Roman, offers a pathway to identify TDEs at $z>4$ and probe black hole demographics at early cosmic times.
@@ 1 Introduction
A tidal disruption event (TDE) occurs when a star passes close enough to a supermassive black hole (SMBH) that the tidal forces on the star overcome its self-gravity, and destroy the star (Hills, 1975; Rees, 1988; Evans & Kochanek, 1989; Ulmer, 1999). This process leads to a luminous flare ($\sim 10^{43}$erg/s), as approximately half of the mass of the star is accreted and the other half is ejected at high velocity (Strubbe & Quataert, 2009; Gezari et al., 2012). This highly transient accretion event contrasts with the more long term, sustained accretion that is associated with active galactic nuclei (AGN) (Begelman et al., 1984; Krolik, 1999; Event Horizon Telescope Collaboration et al., 2019). Despite decades of dedicated efforts, several outstanding questions surrounding the formation and evolution of SMBHs persist. In particular, the uncertain origin of the first SMBHs has outlined the need for robustly measuring the early Universe SMBH mass distribution (Volonteri et al., 2003; Begelman et al., 2006). TDEs are one of the only phenomena, alongside AGN, that present a luminous observable very close to a SMBH. Unlike AGN, which prefer very massive SMBH hosts (Kauffmann et al., 2003), TDEs of main-sequence stars by non-rotating (Schwarzschild) black holes can only occur for lower mass SMBHs because the disrupting black hole has an upper mass limit of $M_{BH}<10^{8}M_{\odot}$ (Hills, 1975). Beyond this mass, the tidal radius $R_{t}=R_{\star}(M_{BH}/M_{\star})^{1/3}$ is within the Schwarzschild radius $R_{s}=2GM/c^{2}$. The TDE rate’s sensitivity to the black hole mass make TDEs an excellent probe of the SMBH mass function (Stone & Metzger, 2016; Kochanek, 2016). The TDE rate is also affected by the stellar mass function, so a redshift-evolving TDE rate could also be sensitive to the initial mass function (IMF) (Magorrian & Tremaine, 1999). Additionally, stellar dynamics calculations imply that galaxies with lower-mass SMBHs have higher intrinsic TDE rates (Stone & Metzger, 2016). Therefore, TDEs are ideal probes of distant, lower-mass SMBHs and the stars which surround them.
@@
Typically TDEs are discovered through wide-field, time-domain optical surveys (e.g. the Zwicky Transient Facility (ZTF) Graham et al., 2019; Bellm et al., 2019). They are slow (lasting 100100200200 days), blue (peaking in the UV) transients that do not typically cool as they fade (Gezari, 2021). TDEs last on scales of months, typically fading from peak in $\sim 200$ days (van Velzen et al., 2021). The origin of the early-time UV/optical emission is debated, and still not fully understood (Ryu et al., 2023; Dai et al., 2018; Roth et al., 2020). In this work, we adopt a fully empirical, observation-driven model of TDE spectra and light curves.
@@
Previously, Kochanek (2016) has used semi-empirical modeling of TDE rates to predict intrinsic and observed rates of TDEs as a function of redshift. They predict that, due to the observed “downsizing” of AGN mass over redshift (as modeled in Shankar et al., 2009), the intrinsic TDE rate will decrease by a factor of $5$ by $z=1$. As a result, they assert that the observed rate of high-redshift TDEs, given ground-based wide-field time domain surveys (focusing on Pan-STARRS1 Chambers et al., 2016), will be even lower. Here, we extend to high redshifts by leveraging deep space-based imaging without the use of time series. We assert in a companion work Karmen & et al. (in prep.) that, given the observed evolution of galaxy and SMBH properties with redshift, the TDE rate at high redshifts is likely greatly enhanced. As a result, in upcoming surveys such as the Roman Space Telescope High Latitude Time Domain Survey (Rose et al., 2021) and the Rubin Observatory’s Legacy Survey of Space and Time (LSST Ivezić et al., 2019) it may be possible to study populations of TDEs at high redshifts. Inayoshi et al. (2023) predicts the properties of high-redshift TDEs embedded within gas-rich AGN environments, in which the TDE occurrence rates are enhanced due to their compact and dense environments, compared to typical TDEs which occur in post-starburst galaxies. In this work, we complement this by a search exclusively for TDEs in galaxies not containing luminous AGN in order to utilize observed properties from low-redshift TDE searches, which exclude all AGN-hosting galaxies.
@@
In this work, we report the discovery of a $z=5.02^{+1.32}_{-1.11}$ TDE candidate in the COSMOS-Web survey (Casey et al., 2022). In Section 2, we present a new methodology for identifying TDE candidates in deep imaging using colors and morphology. We compare to potential contaminants, and pose methods of confirming candidate TDEs. In Section 3 we apply this methodology to the COSMOS-Web survey data, and use historic imaging of the COSMOS field to evaluate the potential variability of objects we select. We present the properties of HZTDE-1, a transient high-redshift TDE candidate in Section 4. We compare the TDE model to potential supernova models for this object. We discuss the need for spectroscopy of this object, and its implications if validated in Section 5. We summarize our work in Section 6.
2 Identifying High-Redshift TDEs in Deep Surveys
In this section, we model the signatures of TDEs similar to the local population, and develop a method for identifying them in wide field infrared surveys using the colors and morphologies. To first order, the UV/optical SED of a TDE is well-approximated by a black body with a typical temperature of $\sim 10^{4.3\pm 0.15}~K$ (van Velzen et al., 2020, 2021; Hammerstein et al., 2023; Yao et al., 2023). We simulate TDEs with the the same temperature distribution as the latest sample of spectroscopically confirmed TDEs from ZTF (Yao et al., 2023), and simulate their volumetric density by integrating over the double-power law fit to the ZTF $g$-band luminosity function (their Equation 16). TDEs typically do not cool as they fade, so we assume constant-temperature blackbodies for their entire evolution. Their light curves are observationally well-parameterized in van Velzen et al. (2021) as follows:
$L_{\nu}(t)$ $=L_{\nu_{0}\,\rm peak}~\frac{B_{\nu}(T_{0})}{B_{\nu_{0}}(T_{0})}$
@@ 2 Identifying High-Redshift TDEs in Deep Surveys
with the luminosity in the $g$-band in erg s-1 $42.68<\log(L_{g})<44.68$, the Gaussian rise time $0.4<\log(\sigma)<1.3$ days and the exponential decay time $1.2<\log(\tau)<2.3$ days from ZTF. When simulating TDEs, we step over redshift in shells of size $\Delta z=0.2$. At a given $z$, we sample over their temperatures and luminosities in a grid starting with the faintest TDE at a given redshift that is visible given the survey limits. We sample their luminosities in step sizes inversely proportional to their probability densities, giving more dense sampling to the higher-probability luminosities. At each redshift and luminosity, we simulate a TDE with every temperature observed in the Yao et al. (2023) ZTF TDE sample. At each given luminosity, we allow TDEs to fade as $L\propto t^{-5/3}$ until they reach the magnitude limit of the filter/survey observing them, to account for the time evolution of the TDE.
@@ 2.1 The Host model
Figure 1: Apparent magnitudes of the median TDE from our simulations (solid lines) compared to the median ZTF host galaxy (dashed lines Hammerstein et al., 2021). The shaded regions represent the 5th and 95th percentile for the TDE hosts. 95% of the hosts drop out by $z\sim 2.2$ in F115W, and by $z\sim 4.1$ in F [^4] W.
@@
We base our predictions of properties of TDE host galaxies on local observations of TDE hosts. It has been observed that TDE hosts are disproportionately in the green valley with compact nuclei (Hammerstein et al., 2021), and may show a preference for post-starburst galaxies (French et al., 2016). Because of this, we base our host galaxy models on the total sample of 30 ZTF TDE hosts from (Hammerstein et al., 2021) to capture their observed properties. To extend the observed host spectra to their rest-frame UV, we use Prospector to model their spectra and perform synthetic photometry (Johnson et al., 2021). We use the parameters from the original fits of the ZTF TDE hosts in Hammerstein et al. (2021), which are estimated by maximum a posteriori fits of Flexible Stellar Population Synthesis models (Conroy et al., 2009) to broadband photometry from SDSS, Pan-STARRS, and GALEX. When calculating properties of TDE hosts, we simulate every TDE host from ZTF to account for their full diversity of properties. We find that, due to the k-correction and the homogeneous colors of the TDE hosts, they typically are not visible beyond $z=3.5$ as seen in Figure 1.
@@
We also use these ZTF TDE hosts as a basis for the physical sizes for typical TDE hosts. We query SDSS photometry for the TDE hosts in ZTF, and convert their angular sizes to physical sizes (Blanton et al., 2017; Hammerstein et al., 2023). We find that ZTF TDE host galaxies contain 90% of petrosian flux in a diameter of $8\leq D_{90}\leq 20$ kiloparsecs. This corresponds to range of angular diameters between $1.4$ and $4.7$ arcsec at $z=8$. This angular size will increase at even higher redshifts due to the decrease in angular diameter distance with increasing redshift beyond $z\sim 1.5$ (Hogg, 1999). As NIRCam has a resolution of 0.07 arcsec at 2 microns, all TDE host galaxies above the survey magnitude limit will be resolvable as an extended source in COSMOS-Web. At early epochs ($z\sim 7$), galaxies are expected to be intrinsically smaller (Grazian et al., 2012; Kawamata et al., 2018), such that even TDEs in unusually bright or blue hosts will not have spatially extended emission.
@@
Because all of the host galaxies we observe are at $z<1$, the galaxies we use are of higher mass and more luminous than those which we expect at high redshifts. High redshift galaxies are expected to be, and now observed to be, primarily lower-mass, blue, star-forming galaxies. While there is a minority of these galaxies which are exceptionally luminous without SMBHs above the Hills mass due to extremely active star formation (e.g. Franco et al., 2024; Casey et al., 2024; Castellano et al., 2024), it would be extremely difficult to detect a UV/optical TDE in one because its star formation is bright and blue, making the TDE emission a sub-dominant component of its UV SED. Therefore, we do not consider them in our analysis. However, a population of TDEs which occur primarily in star-forming galaxies have been observed to be infrared-bright with no UV/optical component to their emission (Masterson et al., 2024), making their rates relevant to wide-field surveys using MIRI. For our search in COSMOS-Web, the local TDE hosts we observe are effectively upper limits for higher-redshift host galaxy luminosity. In summary, TDEs trace lower-mass SMBHs which trace, in turn, lower-mass galaxies. These are expected to be even less massive at high redshifts, therefore intrinsically fainter, and are k-corrected into the mid-infrared. We therefore expect the vast majority of TDEs at $z>4$ to appear hostless.
@@ 2.3 High Redshift TDEs
We find that, given our simulated TDEs and assumptions about their typical host galaxies, most high-redshift ($z>4$) TDEs will be point sources with unique colors. At $z>4$, a typical TDE host galaxy drops in apparent magnitude below the detection limits of the COSMOS-Web survey, as seen in Figure 2. A TDE near peak brightness is visible to beyond $z\sim 10$. The SED of a high-redshift TDE is dominated by its blackbody emission, which is redshifted from the rest-frame UV to the observer-frame near-infrared NIR. Because a redshifted blackbody is another blackbody, there is a degeneracy in color-color space between temperature and redshift, meaning all idealized TDEs lie upon a nearly straight line in 2D color-color space (slightly curved due to relativistic Doppler shift and dust extinction). In our selection for high-redshift TDEs, we restrict to a region that follows the color-color relationship in any given color combination. This is shown as the three dark purple lines outlining the TDE region in Figure 4. We also fit bounds to each individual color, which, together with the parabola fits, are presented in Table 1. We fit a parabola to the color-color relationship of TDEs in each filter combination through maximum likelihood optimization with the assumption of uniform uncertainties, which is mathematically equivalent to least squares fitting. Each curve is parameterized as:
@@
where $C_{1}$ and $C_{2}$ are two colors. The values of coefficients $a_{1}$, $a_{2}$, and $a_{3}$ for each pair of colors are listed in Table 1, alongside the bounds on $C_{1}$ and $C_{2}$ form our simulations. We place bounds on the redshift range we simulate, on the low end by the redshift beyond which TDEs will be point sources, and on the high end by the redshift at which Lyman absorption renders them undetectable. Additionally, we fit the minimum apparent magnitude expected for a TDE at any redshift of a given color, shown in Figure 4, as
@@
To check our blackbody idealization, we take the following TDEs that have Hubble UV spectroscopy, and simulate them at high redshifts: iPTF16fnl, ASASSN-15oi, AT2022dsb, AT2018bsi, and ASASSN-14li (Charalampopoulos et al., 2022; Cenko et al., 2016; Blagorodnova et al., 2019; Brown et al., 2018; Hung et al., 2019; Hoogendam et al., 2024). The spectra used to create these templates can be accessed via doi: 10.17909/9m1b-2s28. We extend the UV spectrum into the optical (>2600>2600 Å) with a blackbody with the temperature fit by Mummery & van Velzen (2024), redshift the TDE spectra, and add in Lyman absorption shortward of rest-frame 1216Å (Inoue et al., 2014). We show these templates in Figure 3. These TDEs have colors centered on the relationship predicted, and scatter due to emission and absorption features, as well as continuum excess farther into the UV. This scatter, typically $\sim 0.2$ mag, defines the region near our idealized color-color relationships in which we expect actual TDEs to exist (the shaded region on the figure).
@@
Figure 3: Example templates used to test the TDE color-magnitude cuts, shifted to the rest frame. The red line is the HST spectrum, purple dotted is the blackbody used to extend the spectrum, and photometry is in blue. We switch from the HST spectrum to the blackbody at rest-frame wavelength $2600$ Å, which is shown in the changing opacity of each line. We simulate this spectrum at higher redshifts to test our color cuts.